OpenStudy (anonymous):

What is b^2/5 in radical form?

OpenStudy (deepolisnoob):

Is it:\[b ^{2/5}\] or \[(b^{2})/5\]

OpenStudy (anonymous):

the first one

OpenStudy (deepolisnoob):

You have \[b ^{n/m}\]. You find your radical for by placing your m value on the root of the equation, so you should have \[\sqrt[m]{b^n}\] If you plug your equation into that formula, you should get: \[\sqrt[5]{b^2}\]

OpenStudy (anonymous):

\[\sqrt[5]{b}^2\]

OpenStudy (anonymous):

Okay thank you that helps a lot

OpenStudy (deepolisnoob):

Anytime :) \[\sqrt[5]{b^2}\] and \[(\sqrt[5]{b})^2\] all have the same value; any of the 2 would be correct